sir,how focallength of a lens changes when red light is replaced by blue light??????????


 

Grade:11

1 Answers

Swapnil Saxena
12 years ago

        According to Cauchy equation

 n(\lambda) = A + \frac {B}{\lambda^2} + \frac{C}{\lambda^4} + \cdots,
Hence the refractive  index of a substance is inversely proptional to the wavelength.

\frac{1}{f} = (n-1) \left[ \frac{1}{R_1} - \frac{1}{R_2} + \frac{(n-1)d}{n R_1 R_2} \right],

Since according the Lens makers formula, the refractive index of a material is directly dependent on its refractive index. Hence in this way we can say that the focal length is dependent on the wavelength of light . Hence blue and red light will show a great variation in thier focal lengths.

In fact his is the cause beind the chromatic abberation produced by the glasses.